OpenStudy (anonymous):

Given the function f(x) = -3x3 + 9x2 - 2x + 3, what part of the function indicates that the left end starts at the top of the graph?

OpenStudy (lucaz):

the highest degree is 3, so for the left end(the negative values) you always get (-x)^3=-number, and the coefficien is also a negative number, -3(x)^3, so you end up with (-)*(-)=(+), for any negative value of x you will get a positive value

OpenStudy (anonymous):

oh okay thank you

OpenStudy (lucaz):

no problem

OpenStudy (anonymous):

can i ask for another problem

OpenStudy (anonymous):

@lucaz can i get help on another problem?

OpenStudy (lucaz):

what is it?

OpenStudy (anonymous):

Calculate the average rate of change for the function f(x) = x4 + 3x3 - 5x2 + 2x - 2, from x = -1 to x = 1.

OpenStudy (lucaz):